OpenStudy (anonymous):

you have a rectangular area with a perimeter of 40ft. What is the maximum area?

OpenStudy (tkhunny):

What is the equation of the perimeter, given length = L and width = W?

OpenStudy (anonymous):

all the question states is the fixed perimeter of 40 ft

OpenStudy (tkhunny):

You didn't answer my question. Answer it, please.

OpenStudy (anonymous):

im guessing you use A=p(1/2)-1^2

OpenStudy (tkhunny):

Really? How about 2w + 2l = 40 It's just the perimeter.

OpenStudy (babalooo):

yes, we know the perimeter is 40 ft. but what makes up the perimeter of a rectangle. or more simply, given a rectangle, how is the perimeter calculated. just basic understanding to make sure you understand the terms.

OpenStudy (babalooo):

lol ok @tkhunny kinda gave it away. so next question, what is the area equation of a rectangle

OpenStudy (anonymous):

the out side of the rectangle...

OpenStudy (babalooo):

what is the formula for area of a rectangle?

OpenStudy (anonymous):

A=WL

OpenStudy (babalooo):

and what's the formula for perimeter of a rectangle?

OpenStudy (anonymous):

P=2l+2w

OpenStudy (babalooo):

yes and we know P=40 right? so now you just have to find 2 values l and w that fulfills the perimeter equation that gives the biggest area using the area formula

OpenStudy (anonymous):

i got all that. i just dont know how to find the l in the formula. you end up getting A=20-l^2

OpenStudy (babalooo):

how do you get that A=20-l^2

OpenStudy (anonymous):

plug it into A=p(1/2)-l^2 p=40

OpenStudy (anonymous):

with the other formula you get 40=2l+2w. and we dont know what l or w is.

OpenStudy (babalooo):

sorry, i don't quite recognize your A=p(1/2)-l^2 formula.. what's the formula called.

OpenStudy (babalooo):

can't you just trial go through trial and error with values of l and w to get max area?

OpenStudy (anonymous):

its one of the formulas given in the text book. its the first time ive seen it before also

OpenStudy (babalooo):

i found this, think it'll help you https://ca.answers.yahoo.com/question/index?qid=20061220071413AAxAFaU

OpenStudy (tkhunny):

What is \(20 - l^{2}\)? That doesn't mean anything. \(Perimeter(L,W) = 2*L + 2*W = 40\) \(Area(L,W) = L*W\) From the perimeter, we have \(L = (40 - 2W)/2 = 20 - W\) Substituting in to the Area definition, \(Area(L) = L*(20-L) = 20L - L^{2}\) \(Area(L) = 0\;for\;L = 0\;and\;L = 20\). This puts the maximum at \(L = 10\) which implies \(W = 10\).

OpenStudy (anonymous):

thanks for the help now that i see it it makes sense

OpenStudy (tkhunny):

I do wish I could have seen some of YOUR work? Thsi wen ton too long and got too confusing. I had to just spill it.

OpenStudy (anonymous):

lol i know what your saying. i haven't seen this stuff in 8 years